OpenStudy (anonymous):

what is the simplified solution to 3 7/8 - 1 3/4

OpenStudy (anonymous):

what is 3 -1 ?

OpenStudy (anonymous):

2

OpenStudy (anonymous):

right, it's 2 :) so we already have 2 ...

OpenStudy (anonymous):

fractions: 7/8 - 3/4 can we already calculate the result like this?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

it is only allowed to subtract/add fractions with the same denominator! /8 and /4 is not the same denominator so we can't yet add/subtract those fractions

OpenStudy (anonymous):

if you multiply numerator and denominator of a fraction with the same number, the fraction remains the same: 1/2 = 2/4 = 3/6. 3/6 has higher numbers, but 3 of 6 is also a half

OpenStudy (anonymous):

ok?

OpenStudy (anonymous):

you can e.g. "extend" 1/2 by 5 like this: 5 / 10 <- this is also a half

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

7/8 and 3/4 is it possible to "extend" one of the fractions and have as result two fractions with the same denominator? (remember - same denominator is required to add/subtract fractions)

OpenStudy (anonymous):

do you know how to extend a fraction?

OpenStudy (anonymous):

noo

OpenStudy (anonymous):

ok :) you multiply numerator AND denominator with some number before: \[\frac{ 1 }{ 2 }\]after extending with 5:\[\frac{ 1 \times 5 }{ 2 \times 5 } = \frac{ 5 }{ 10 }\]

OpenStudy (anonymous):

oh i see

OpenStudy (anonymous):

\[\frac{ 7 }{ 8 } ,...\frac{ 3 }{ 4 }\]we want to subtract 3/4 from 7/8, but subtraction/addition is only allowed when fractions have the same DENOMINATOR the denominator is the number under the line, is it the same here?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

correct is it possible to extend one of the fractions so they will have the same denominator though? which one and extend by what number?

OpenStudy (anonymous):

multiply them

OpenStudy (anonymous):

\[\frac{ 7 }{ 8 } ,...\frac{ 3 }{ 4 }\]They need the same number under the line, how can that be accomplished? By extending one of the fractions. Which one and by what?

OpenStudy (anonymous):

i have no idea

OpenStudy (anonymous):

\[\frac{ 7 }{ 8 } ,...\frac{ 3 }{ 4 }\]you can extend the second fraction by some number to make it a /8 also ;)

OpenStudy (anonymous):

by what number should the second fraction, 3/4, be extended? so that its denominator will be an 8 ?

OpenStudy (anonymous):

the denominator must become an 8 because 7/8 has the denominator 8 extend 3/4 by what to accomplish this ?